# CrossSDF: 3D Reconstruction of Thin Structures From Cross-Sections

### v1

@@ 3.3 Hybrid Encoding
Optimizing a neural field directly on 3D coordinates can result in over-smooth geometry, as neural networks are biased towards low-frequency functions on low dimensional domains [38]. As a result, positionally encoding 3D input coordinates with Fourier features has become common practice to learn detailed implicit fields [15]. Grid-based encodings offer an alternative, encoding positions by interpolating learned feature vectors stored on grid vertices [37, 24, 22]. Recently, the use of spatial hashing has allowed grid-based encodings to scale to higher resolutions, and therefore capture even finer surface details [29, 22]. Inspired by this, we employ a hash-grid encoding to capture thin structures. However, these encodings are not directly suitable to our problem, since they introduce grid artifacts visible in the resulting geometry. These artifacts are due to linear interpolation that is performed across grid cell boundaries, leading to discontinuities that manifest as sharp creases or ridges in the learned surface (see Fig. 8). This effect is especially pronounced for our task, since the sparsity of the cross-sections means that large 3D regions between input planes are not intersected and we rely exclusively on priors for interpolation. To mitigate these issues, we introduce an alternative encoding strategy combining random Fourier features [38] with multi-resolution hash-grid features [29]. We define the random Fourier feature (RFF) encoding $\gamma_{\text{RF}}:\mathbb{R}^{3}\to\mathbb{R}^{d}$ as follows:
@@ 4.1 Results
Figure 6: Qualitative comparison of hepatic vessel reconstruction slices derived from a real CT scan.
@@
We also test CrossSDF on real word data from the IRCADb-01 [36] and medical decathlon [4] datasets, which contains anonymized and annotated CT scans in DICOM format. Qualitative results are shown in Fig. 6 on a hepatic vessel with 61 slices and 5 millimeters between slices, which in medical imaging is considered sparse. We report more qualitative and quantitative results in the Appendix.
@@ A.1 CT Scan Results
We present qualitative reconstructions of real anatomical structures with two vessel structures from the IRCADb-01 dataset [36] and two from the Medical Decathlon dataset [4], using pixel-level human-annotated segmentations. We evaluate CrossSDF on hepatic vessels from slices spaced 5mm apart, and liver vessels at 1mm apart. We display the results across different methods in Fig. A1 which demonstrates our model’s ability to handle clinically relevant challenges. As no ground truth 3D geometry is available, we created training and test splits by withholding $\approx 10\%$ of the slices and reported the 2D intersection-over-union (IoU) on the test set. To do so, we use integer division to compute a slice skipping frequency to withhold cross-sections. For example, for the hepatic vessel of patient 31, we remove every $61//10=6^{th}$ cross-section to compute 2D IoU. The quantitative results can be found in Tab. A1.
@@
Method Hepatic Vessel Liver
Patient 31 Patient 33 Patient 13 Patient 14
(61 Slices) (71 Slices) (149 Slices) (75 Slices)
CrossSDF (Ours) 0.79 0.63 0.83 0.88
OReX [33] 0.71 0.09 0.77 0.81
Poisson [19] 0.24 0.55 0.21 0.84
Table A1: Results on real CT scans. We report the average 2D IoU on a held out set of evaluation slices.
@@ A.2 Additional Synthetic Results
We display further qualitative comparisons in Figs. A2 and A4. In Fig. A3 we also show results from Bermano et al. [9], a non-neural method that can handle arbitrary cross-section orientations, and works by a barycentric blending of indicator functions in the cells of the planar arrangements of the cross-sections. As can be seen in the figure, their approach is susceptible to laddering artifacts, and fails to run on more complex scenes due to poor scalability.
@@
In addition to the baselines reported in the main paper, we conduct a further ablation by removing all of the key components of our CrossSDF approach; i.e., the adaptive encoding, Fourier features, and symmetric difference loss. We denote this as the ‘Standard Hash-grid’ baseline. The quantitative results for the Elephant scene are presented in Table Tab. A2. Removing these components leads to a degradation in performance, with the full model achieving the lowest CD and HD values.
@@ A.3 Supplementary Video
In the attached supplementary video we provide comparison of a reconstruction from our approach compared to the baselines. We also visualize the optimization process on a different scene.
@@ Appendix B Implementation Details
We implemented our solution in PyTorch and used the accelerated tiny-cuda-nn [28] implementation for our hash-encoding module. All scenes were trained using a single 24GB NVIDIA RTX 4090. We train our network for a total of 500 epochs. The running time is commensurate with the complexity of the scene, taking as little as 2020~{}2020 minutes in the simplest cases (e.g., Elephant in Fig. A4) to 55~{}55 hours in the most complex scenes (e.g., Alveolis in Fig. A2). Our method does not have to create a (cubic-complexity) arrangement of cross-sectional planes, unlike [9, 46], and thus we scale better with the input complexity, only depending on the number of samples. Our network hyper-parameters are not tuned for each dataset. We use the same weights for losses and hash encoding resolution for all the datasets displayed. The mesh is extracted using marching cubes at $512^{3}$ at the end of training.

### v3

@@ 1 Introduction
22footnotetext: Project website: https://iamsalvatore.io/cross_sdf
@@ 3.3 Hybrid Encoding
Optimizing a neural field directly on 3D coordinates can result in over-smooth geometry, as neural networks are biased towards low-frequency functions on low dimensional domains [38]. As a result, positionally encoding 3D input coordinates with Fourier features has become common practice to learn detailed implicit fields [15]. Grid-based encodings offer an alternative, encoding positions by interpolating learned feature vectors stored on grid vertices [37, 24, 22]. Recently, the use of spatial hashing has allowed grid-based encodings to scale to higher resolutions, and therefore capture even finer surface details [29, 22]. Inspired by this, we employ a hash-grid encoding to capture thin structures. However, these encodings are not directly suitable to our problem, since they introduce grid artifacts visible in the resulting geometry. These artifacts are due to linear interpolation that is performed across grid cell boundaries, leading to discontinuities that manifest as sharp creases or ridges in the learned surface (see Fig. 8). This effect is especially pronounced for our task, since the sparsity of the cross-sections means that large 3D regions between input planes are not intersected and we rely exclusively on priors for interpolation. To mitigate these issues, we introduce an alternative encoding strategy combining random Fourier features (RFFs) [38] with multi-resolution hash-grid features [29]. We define the random RFF encoding $\gamma_{\text{RF}}:\mathbb{R}^{3}\to\mathbb{R}^{d}$ as follows:
@@ 4 Experiments
We also test on real word data from the IRCADb-01 [36] and medical decathlon [4] datasets, which contains anonymized and annotated CT scans in DICOM format. As no ground truth 3D geometry is available, we created training and test splits by withholding $\approx 10\%$ of the slices.
@@ 4.1 Results
Figure 6: Qualitative comparison of hepatic vessel reconstruction (partial view shown), with 61 slices derived from a real CT scan.
Method Hepatic Vessel Liver
Patient 31 Patient 33 Patient 13 Patient 14
(61 Slices) (71 Slices) (149 Slices) (75 Slices)
CrossSDF (Ours) 0.79 0.63 0.83 0.88
OReX [33] 0.71 0.09 0.77 0.81
Poisson [19] 0.24 0.55 0.21 0.84
Table 3: Results on real CT scans. We report the average 2D IoU on a held out set of evaluation slices.
@@
CT Scans.
CrossSDF displays superior performance on real-world CT scan data, see qualitative example in Fig. 6, and quantitative results in Tab. 3. Please see the Appendix for more results.
@@ A.1 CT Scan Results
We present qualitative reconstructions of real anatomical structures with two vessel structures from the IRCADb-01 dataset [36] and two from the Medical Decathlon dataset [4], using pixel-level human-annotated segmentations. We evaluate CrossSDF on hepatic vessels from slices spaced 5mm apart, and liver vessels at 1mm apart. We display the results across different methods in Fig. A1 which demonstrates our model’s ability to handle clinically relevant challenges. As no ground truth 3D geometry is available, we created training and test splits by withholding $\approx 10\%$ of the slices and reported the 2D intersection-over-union (IoU) on the test set. To do so, we use integer division to compute a slice skipping frequency to withhold cross-sections. For example, for the hepatic vessel of patient 31, we remove every $61//10=6^{th}$ cross-section to compute 2D IoU. The quantitative results can be found in Tab. 3.
@@
Thin Structures
IoU $\uparrow$ Neural-IMLS [40] 0.050 0.048 0.44 0.045 0.029 0.040 0.71 0.74 0.033 0.021 0.056 0.045
Screened Poisson [19] 0.84 0.72 0.85 0.79 0.78 0.68 0.85 0.89 0.28 0.71 0.89 0.86
OReX [33] 0.37 0.12 0.64 0.33 0.51 0.63 0.96 0.97 0.32 0.42 0.59 0.74
CrossSDF (Ours) 0.87 0.76 0.94 0.90 0.88 0.78 0.99 0.97 0.77 0.70 0.96 0.94
Table A1: Volume IoU comparison across thin structures.
@@ A.2 Additional Synthetic Results
We display further qualitative comparisons in Figs. A2 and A4. We also report quantitative comparisons of 3D Volume IoU in Tab. A1. In Fig. A3 we also show results from Bermano et al. [9], a non-neural method that can handle arbitrary cross-section orientations, and works by a barycentric blending of indicator functions in the cells of the planar arrangements of the cross-sections. As can be seen in the figure, their approach is susceptible to laddering artifacts, and fails to run on more complex scenes due to poor scalability.
@@
In addition to the baselines reported in the main paper, we conduct a further ablation by removing all of the key components of our CrossSDF approach; i.e., the adaptive encoding, Fourier features, and symmetric difference loss. We denote this as the ‘Standard Hash-grid’ baseline. The quantitative results for the Elephant scene are presented in Tab. A2. Removing these components leads to a degradation in performance, with the full model achieving the lowest CD and HD values.
@@ A.3 Supplementary Video
In the video, available on our website, we provide comparison of a reconstruction from our approach compared to the baselines. We also visualize the optimization process on a different scene.
@@ Appendix B Implementation Details
We implemented our solution in PyTorch and used the accelerated tiny-cuda-nn [28] implementation for our hash-encoding module. All scenes were trained using a single 24GB NVIDIA RTX 4090. We train our network for a total of 500 epochs. The running time is commensurate with the complexity of the scene, taking as little as 2020~{}2020 minutes in the simplest cases (e.g., Elephant in Fig. A4) to 55~{}55 hours in the most complex scenes (e.g., Alveolis in Fig. A2). Our method does not have to create a (cubic-complexity) arrangement of cross-sectional planes, unlike [9, 46], and thus we scale better with the input complexity, only depending on the number of samples. Our network hyper-parameters are not tuned for each dataset. We use the same weights for losses and hash encoding resolution for all the datasets displayed. The mesh is extracted using marching cubes at $512^{3}$ at the end of training (similarly for competing methods).